# Composing Agents to Minimize Worst-case Risk

### v1

@@ Abstract
From software development to robot control, modern agentic systems decompose complex objectives into a sequence of subtasks and choose a set of specialized AI agents to complete them. We formalize an agentic workflow as a directed acyclic graph, called an agent graph, where edges represent AI agents and paths correspond to feasible compositions of agents. When deploying these systems in the real world, we need to choose compositions of agents that not only maximize the task success, but also minimize risk where the risk captures requirements like safety, fairness, and privacy. This additionally requires carefully analyzing the low-probability (tail) behaviors of compositions of agents. In this work, we consider worst-case risk minimization over the set of feasible agent compositions. We define worst-case risk as the tail quantile—also known as value-at-risk—of the loss distribution of the agent composition where the loss quantifies the risk associated with agent behaviors. We introduce an efficient algorithm that traverses the agent graph and finds a near-optimal composition of agents by approximating the value-at-risk via a union bound and dynamic programming. Furthermore, we prove that the approximation is near-optimal asymptotically for a broad class of practical loss functions. To evaluate our framework, we consider a suite of video game-like control benchmarks that require composing several agents trained with reinforcement learning and demonstrate our algorithm’s effectiveness in approximating the value-at-risk and identifying the optimal agent composition.
@@ 1 Introduction
Modern agentic workflows orchestrate specialized AI agents to tackle complex tasks demanding a diverse range of skills. Typically, a high-level planning agent decomposes the main task into a sequence of subtasks and selects appropriate worker agents to execute them. These worker agents can be powerful pre-trained models, such as large language models (LLMs) and vision-language models (VLMs), or control policies trained with reinforcement learning or imitation learning. Such systems have demonstrated considerable success in automating tasks across various domains, including software development [34, 24, 14] and robot control [16, 9, 33].
In addition to maximizing task success, deploying agentic systems in the real world requires minimizing various forms of risk. These risks encompass requirements like safety, fairness, and privacy. The stochastic nature of both environment and agents means that we have to rigorously analyze low-probability but high-consequence tail behaviors that emerge when multiple agents interact in sequence. Thus, quantifying and minimizing these compositional risks becomes essential for responsible deployment.
@@
Figure 1: MouseNav environment (left) and its agent graph (right).
We introduce our framework through an illustrative example. Consider the toy environment MouseNav shown in Figure 1 where we are tasked with navigating a robot mouse to the slice of cheese while avoiding the traps. Suppose we have access to four control agents: $\pi_{1},\pi_{2}$ that can take the mouse from the initial room $M$ to rooms $A$ or $B$, and $\pi_{3},\pi_{4}$ that can navigate from rooms $A$ or $B$ to the cheese room $C$. This gives us two compositions of agents to pick from: $M\xrightarrow{\pi_{1}}A\xrightarrow{\pi_{3}}C$ which takes the mouse through room A (green path) and $M\xrightarrow{\pi_{2}}B\xrightarrow{\pi_{4}}C$ which takes it through room B (blue path). In this work, we represent feasible agent compositions as an agent graph—a directed acyclic graph in which each agent is represented by an edge and an agent composition is given by a directed path in the graph (as seen in Figure 1). Next, we quantify the risk associated with an agent’s behavior with real-valued loss functions. For example, we can quantify the safety of the control agent $\pi_{1}$ with a function $L_{1}$ that takes a trajectory of motion of the mouse and returns the negative minimum distance between the trajectory and the traps. Note that we define the loss to be the negative minimum distance because we prefer being as far as possible from the traps. Subsequently, we define the loss of a composition of agents as the maximum of the losses incurred by each agent since we are concerned with providing worst-case guarantees. In the case of the composition $M\xrightarrow{\pi_{1}}A\xrightarrow{\pi_{3}}C$, the loss would be given by $\max(L_{1},L_{3})$ and again corresponds to the negative minimum distance between a trajectory and the traps except now the trajectory is a sequence of two trajectories: the first one from executing $\pi_{1}$ and then from $\pi_{3}$.
@@
Given an agent graph, our objective is to find the path representing an agent composition that minimizes the worst-case risk. Worst-case risk is defined as the $(1-\alpha)$-quantile of the loss distribution of the agent composition for a user-defined risk budget $\alpha>0$. This is also commonly known as the value-at-risk at level $\alpha$ denoted $\textnormal{VaR}_{\alpha}$. Rephrasing in the context of the MouseNav example, minimizing the worst-case risk would maximize a constant $D>0$ such that the chosen composition of agents ensures that the robot mouse stays at least distance $D$ away from the traps with probability at least $1-\alpha$. In the rest of the work, we only assume black-box sampling access to the agents and the input distributions, and so we have to estimate the value-at-risk using samples. A straightforward approach to minimizing the worst-case risk is to consider each feasible agent composition, estimate its value-at-risk, and pick the optimal composition. However, this method is inefficient since the graph can have an exponential number of compositions as a function of the number of agents.
@@
To overcome this hurdle, we propose an efficient algorithm that only depends polynomially on the number of agents. It works by incrementally building a near-optimal path while traversing the agent graph. At its heart is a quantile estimation method that approximates the value-at-risk of a composition of agents by dividing the risk budget $\alpha$ amongst the agents and applying a union bound. It then finds the optimal allocation of the risk budget over a discretized set of budgets using a dynamic programming algorithm. We further prove that this approximation is near-optimal asymptotically when the loss of each agent can be represented by an independent random variable. This independence assumption is reasonable in practice especially when the losses are functions of the agent behavior and do not directly impact task accuracy or performance.
@@
We implement our algorithm and test its performance on a series of reinforcement learning (RL) environments. Here, we look at long-horizon tasks that require composing several agents trained with RL to complete them. We consider two forms of risk minimization—safety in navigation and resource consumption—and design suitable loss functions that capture these requirements. We confirm empirically that our algorithm produces tight approximations of the value-at-risk thereby finding optimal agent compositions and also scales favorably to higher number of agents.
@@
    We formalize the worst-case risk minimization objective for agentic workflows as the problem of finding an agent composition that minimizes the value-at-risk of losses quantifying risk (Section 2).
@@
    Then, we introduce an efficient algorithm that approximates this optimization problem with a quantile estimation method using the union bound and dynamic programming. Furthermore, we prove that this approximation is near-optimal asymptotically under independence assumptions. (Section 3)
@@
    Finally, we evaluate the algorithm on compositional reinforcement learning benchmarks and; we show that the algorithm succeeds in finding the optimal path and tightly approximates the value-at-risk (Section 4).
@@ 1.1 Related work
Our work builds upon a large body of literature on automated task planning under uncertainty; we refer to Ghallab et al. [11] for a comprehensive treatment of this subject. Traditionally, the planning problem is modeled as a Markov decision process (MDP) over which the expected (un)discounted cumulative losses are minimized using algorithms like value iteration or reinforcement learning [31]. Towards designing risk-aware agents, recent works have instead considered the optimization of risk measures that look at the tail of the loss distribution instead of the expected loss. Ahmadi et al. [1], Kretínský and Meggendorfer [19] study risk minimization for reachability and mean-payoff objectives in MDPs, while Bastani et al. [3], Wang et al. [32], Greenberg et al. [13] study the same in the context of reinforcement learning. In these works, the considered risk measures fall in the class of coherent risk measures [2] that include measures like the conditional value-at-risk and the entropic value-at-risk. However, all of these works are in contrast to our framework where we aggregate the losses by considering the maximum incurred loss instead of the cumulative loss.
@@
Related lines of research in control and robotics include hierarchical reinforcement learning [17, 7, 35, 15, 21] and task and motion planning (TAMP) [18] in which long-horizon control objectives are decomposed into subtasks that are carried out by motion planners or other low-level controllers trained with RL or imitation learning. Similar frameworks have been proposed for automatic generation of agentic workflows to automate software development tasks [24, 34, 14] that sequence LLM-agents to complete smaller subtasks. These methods typically optimize the overall task success probability and do not consider risk minimization.
Lastly, our work is related to conformal prediction [29, 25] methods for ML models which also requires quantile estimation of the non-conformity score. In particular, conformal prediction has been applied to sequential prediction settings like for compositions of models [27, 20] and for time-series prediction [12, 4, 30, 6]. Li et al. [20], Cleaveland et al. [6] also grapple with the issue of optimal allocation of the error budget along the sequence of predictions when applying the union bound. Li et al. [20] address this using a Bayesian optimization approach while Cleaveland et al. [6] circumvent using the union bound with an alternative approach based on linear complementarity programming. However, these methods only work in the sequential prediction setting and cannot be scaled to a graph of predictions as required in our framework.
2 Worst-case Risk Minimization Over Agent Graphs
@@
In this work, an agent is any machine learning model $f:X\to\mathfrak{D}(T\times Y)$ where $X$ is the input domain, $T$ is the set of computational traces capturing the set of agent behaviors, $Y$ is the output domain, and $\mathfrak{D}(T\times Y)$ is the set of probability distributions over $T\times Y$. For example, for an agent controlling a robot, $X$ is the set of initial states, $T$ is the set of trajectories of the robot in its environment, and $Y$ is the set of final states. In the case of a language model, $X$ is the set of prompts, $Y$ is the set of final responses, and $T$ is the set of traces used to produce the final response. This can be the Chain-of-Thought reasoning traces, or interactions with a tool like retrieval from the web or running code. We consider a distribution over traces and outputs, as opposed to a single output, to enable modeling stochastic agent behaviors along with interactions with a stochastic environment. We only assume sampling access to the distribution over outputs and denote $(t,y)\sim f(x)$ for sampling a trace-output pair $(t,y)$ from the agent $f$ for the input $x$.
@@
We quantify the risk associated with an agent’s trace by a loss function $L:T\to\mathbb{R}$. This formulation allows us to decouple risk, which depends on the agent’s overall behavior, from task accuracy, which only depends on the final output. Thus, the loss function can encode desiderata of the agent’s behavior like safety, privacy, and fairness. For example, in the MouseNav environment from Figure 1, the loss function quantifies how safe a control policy is on a trajectory of the mouse.
@@
Observe that given a distribution over the inputs $\mathcal{D}(X)$, the agent $f$ induces distributions $\mathcal{D}_{f}(T)$ and $\mathcal{D}_{f}(Y)$ over the set of traces and outputs respectively. We can sample $t\sim\mathcal{D}_{f}(T)$ or $y\sim\mathcal{D}_{f}(Y)$ by first sampling $x\sim\mathcal{D}(X)$, then sampling $(t,y)\sim f(x)$, and dropping one of the two components. Evaluating the loss on the distribution of traces gives us a distribution over the losses the agent incurs. Since we are concerned with providing worst-case guarantees, we focus on analyzing tail behaviors of the loss distribution. In particular, the quantity we are interested in optimizing is the value-at-risk at level $\alpha\in(0,1)$, denoted $\textnormal{VaR}_{\alpha}$. $\textnormal{VaR}_{\alpha}$ is a risk measure of the loss distribution that corresponds to the $(1-\alpha)$-quantile of the distribution. In other words, a $(1-\alpha)$ fraction of the traces will have a loss smaller than the $\textnormal{VaR}_{\alpha}$. To formally define it, let $Z_{f}\sim\mathcal{D}_{f}(T)$ be a random variable over the traces of the agent $f$. Then we can express
@@
which is the maximum of losses incurred by the sequence of trajectories $(t_{1},\dots,t_{m})$. This definition is consistent with our goal of providing worst-case guarantees. Nonetheless, there may be other sensible definitions; e.g., when the losses represent monetary cost, a natural quantity is the cumulative loss which has been studied previously in the context of risk minimization [1, 19].
We now have all the ingredients to state our optimization objective over an agent graph that we call the worst-case risk minimization (WoRM) objective.
Objective (WoRM).
Let $G=(V,E,X,T,F,L,s,t,\mathcal{D}_{s})$ be an agent graph and let $\mathcal{P}\subseteq V^{*}$ be the set of directed paths from $s$ to $t$. For a given risk level $\alpha\in(0,1)$ we define the worst-case risk minimization objective as the following optimization problem
@@
3 Finding the path minimizing the worst-case risk
Algorithm BucketedVaR : Efficient algorithm for the WoRM objective
1:$G=(V,E,X,T,F,L,s,t,\mathcal{D}_{s})$: agent graph; $d\in\mathbb{N}_{>0}$: number of buckets; $n$: sample size; risk budget $\alpha\in(0,1)$
2:$B\leftarrow\{0,\frac{\alpha}{d},2\frac{\alpha}{d},\dots,\alpha\}$triangleright Set of risk level buckets
3:VaR:($\textnormal{{VaR}}:(V\times B\to\mathbb{R})\leftarrow$ init empty map $\triangleright$ Map to store VaR estimates

### v2

# Risk-Sensitive Agent Compositions

@@ Abstract
From software development to robot control, modern agentic systems decompose complex objectives into a sequence of subtasks and choose a set of specialized AI agents to complete them. We formalize agentic workflows as directed acyclic graphs, called agent graphs, where edges represent AI agents and paths correspond to feasible compositions of agents. Real-world deployment requires selecting agent compositions that not only maximize task success but also minimize violations of safety, fairness, and privacy requirements which demands a careful analysis of the low-probability (tail) behaviors of compositions of agents. In this work, we consider risk minimization over the set of feasible agent compositions and seek to minimize the value-at-risk of the loss distribution of the agent composition where the loss quantifies violations of these requirements. We introduce an efficient algorithm which traverses the agent graph and finds a near-optimal composition of agents. It uses a dynamic programming approach to approximate the value-at-risk of agent compositions by exploiting a union bound. Furthermore, we prove that the approximation is near-optimal asymptotically for a broad class of practical loss functions. To evaluate our framework, we consider a suite of video game-like control benchmarks that require composing several agents trained with reinforcement learning and demonstrate our algorithm’s effectiveness in approximating the value-at-risk and identifying the optimal agent composition.
@@ 1 Introduction
Modern agentic workflows orchestrate specialized AI agents to tackle complex tasks demanding a diverse range of skills. Typically, a high-level planning agent decomposes the main task into a sequence of subtasks and selects appropriate worker agents to execute them. These worker agents can be powerful generalist models, such as large language models (LLMs) and vision-language models (VLMs), or control policies trained with reinforcement learning or imitation learning. Such systems have demonstrated considerable success in automating tasks across various domains, including software development [33, 25, 15], complex information retrieval [34], scientific discovery [13], and robot control [17, 9, 32, 35].
In addition to maximizing task success, deploying agentic systems in the real world requires minimizing various forms of risk. The stochastic nature of both environment and agents means that we have to rigorously analyze low-probability but high-consequence tail behaviors that emerge when multiple agents interact. Our work follows a rich body of literature in risk-sensitive planning and control that dates back to the 1970s [14, 31] where the objective is to optimize a risk measure of the loss (or reward) distribution as opposed to the expectation. A risk measure is a function that maps a loss distribution to a real number and takes into account attributes like the variance and the tail of the distribution. Some examples include the value-at-risk (VaR) which corresponds to the tail quantile and conditional value-at-risk (CVaR) which is the expected tail loss. In this work, we consider loss functions that quantify violations of requirements such as safety, fairness, and privacy.
In our formalization, we represent an agentic workflow by an agent graph (example in Figure˜1) which is a directed acyclic graph in which edges correspond to agents (denoted $\pi_{1},\dots,\pi_{4}$) and paths correspond to feasible agent compositions that achieve the overall objective. In Figure˜1, we have two compositions of agents to pick from: $p_{A}\coloneqq\pi_{1}\to\pi_{3}$ and $p_{B}\coloneqq\pi_{2}\to\pi_{4}$. Finally, the risk associated with an agent’s behavior is quantified with a real-valued loss function $L_{i}:T_{i}\to\mathbb{R}$ that takes as input a trace of execution of the corresponding agent $\pi_{i}$.
@@
Figure 1: An agent graph
Example 1: DroneNav. Consider four control agents $\pi_{1},\dots,\pi_{4}$ that have been trained to navigate a drone between the rooms $S\xrightarrow{\pi_{1}}A$, $S\xrightarrow{\pi_{2}}B$, $A\xrightarrow{\pi_{3}}F$, and $B\xrightarrow{\pi_{4}}F$. We define loss functions $L_{i}:T_{i}\to\mathbb{R}$ for each agent that capture safety requirements. They take as input the trajectory from a rollout of a policy and return the negative minimum distance between the obstacles in the rooms and the trajectory. Now, consider the task of navigating a drone from room $S$ to the target room $F$ which can be achieved by two paths: one passing through room $A$ and another through room $B$. This can be represented by the agent graph in Figure˜1 where we have the two compositions of agents to choose from. The loss functions of the compositions of agents $p_{A}$ and $p_{B}$ correspond to the negative minimum distance between the obstacles and the sequence of trajectories of the agents in each composition. In particular, for the composition $p_{A}$, the loss function is denoted $L_{p_{A}}:T_{1}\times T_{3}\to\mathbb{R}$ and is given by $L_{p_{A}}(t_{1},t_{3})\coloneqq\max(L_{1}(t_{1}),L_{3}(t_{3}))$ where $t_{1}$ and $t_{3}$ are trajectories from rollouts of $\pi_{1}$ and $\pi_{3}$ respectively. $L_{p_{B}}$ is symmetrically defined.
Example 2: Information retrieval. Suppose we have access to an LLM agent that we are using in an information retrieval and processing task. In this setting, a natural requirement is to minimize the amount of hallucinated information, i.e., data that is not taken from data sources or the LLM context. We define the loss function to be a metric that quantifies the amount of hallucinated information in the outputs which can be a score produced by an LLM-as-a-Judge. Next, consider the task of constructing an agentic system that takes as input the name of a country and produces a graph of the evolution of literacy rates in the country over a period of five decades. Also, suppose we have two data sources—UNESCO and the World Bank—from which the system needs to fetch relevant pages and then write a Python script to generate the plot. This can be represented by the agent graph in Figure˜1 where $\pi_{1}$ and $\pi_{2}$ represent LLM agents tasked with retrieving relevant text data from the two sources, and $\pi_{3}$ and $\pi_{4}$ represent agents that write the script to produce the plot from the data. Yet again, the loss of the composition of agents is given by $L_{p_{A}}(t_{1},t_{3})\coloneqq\max(L_{1}(t_{1}),L_{3}(t_{3}))$ which now corresponds to the maximum amount of information hallucinated by either of the two agents $\pi_{1}$ or $\pi_{3}$ in their corresponding outputs $t_{1}$ and $t_{3}$.
As noted in these examples, when the loss of each agent quantifies violations of safety, fairness, and privacy requirements, the loss of a composition of agents—the quantity we want to minimize—is the maximum of the losses incurred by each agent. In contrast, existing literature in risk-sensitive planning usually considers the cumulative loss [1, 20] which is the natural choice when the losses represent costs.
@@
Given an agent graph, our objective is to find the path representing an agent composition that minimizes the value-at-risk, denoted $\textnormal{VaR}_{\alpha}$, which is the $(1-\alpha)$-quantile of the loss distribution of the agent composition for a user-defined risk budget $\alpha>0$. In other words, minimizing $\textnormal{VaR}_{\alpha}$ minimizes a constant $\ell^{*}>0$ such that with probability at least $1-\alpha$, each agent in the chosen composition of agents incurs a loss of at most $\ell^{*}$. In the rest of this work, we assume only black-box access to the agents and focus on estimating the value-at-risk through sampling.
@@
A straightforward approach to minimizing the value-at-risk is to consider each feasible agent composition, estimate its tail quantile, and pick the optimal composition. However, this method is inefficient since the graph can have an exponential number of compositions as a function of the number of agents. We overcome this hurdle with an efficient algorithm that scales polynomially in the number of agents and incrementally builds a near-optimal path while traversing the agent graph. At its heart is a quantile estimation method that approximates the value-at-risk of a composition of agents by dividing the risk budget $\alpha$ amongst the agents and applying a union bound. It then finds the optimal allocation of the risk budget over a discretized set of budgets using a dynamic programming algorithm. We prove that this approximation is near-optimal asymptotically when the losses of agents are given by independent random variables. This independence assumption holds when the losses are functions of the agent behavior and do not directly impact task accuracy or performance. Indeed, safety, fairness, and privacy are often properties of the behavior and are orthogonal to task success.
@@
We implement our algorithm and test its performance on a series of reinforcement learning (RL) environments. Here, we look at long-horizon tasks that require composing several agents trained with RL to complete them. We consider two tasks—safety in navigation and fair resource consumption—and design suitable loss functions that capture these requirements. We confirm empirically that our algorithm produces tight approximations of the value-at-risk thereby finding optimal agent compositions.
@@
    We formalize the risk minimization objective for agentic workflows as the problem of finding an agent composition that minimizes the value-at-risk of the loss quantifying violations of safety, fairness, and privacy requirements (Section˜2).
@@
    Then, we introduce an efficient algorithm that approximates this optimization problem with a quantile estimation method using the union bound and dynamic programming. Furthermore, we prove that this approximation is near-optimal asymptotically under independence assumptions. (Section˜3)
@@
    Finally, we evaluate our algorithm on compositional reinforcement learning benchmarks using two tasks: safety-critical navigation and fair resource consumption. We design appropriate loss functions for these tasks and demonstrate that our algorithm successfully identifies optimal agent compositions while providing tight approximations of their value-at-risk (Section˜4).
@@ 1.1 Related work
Our work builds upon a large body of literature on automated planning and control under uncertainty [11]. Traditionally, the planning problem is modeled as a Markov decision process (MDP) over which the expected (un)discounted cumulative loss is minimized using algorithms like value iteration or reinforcement learning. Towards designing risk-aware agents, recent works have instead considered the optimization of risk measures that look at the tail of the loss distribution instead of the expected loss. Ahmadi et al. [1], Kretínský and Meggendorfer [20] study risk minimization for reachability and mean-payoff objectives in MDPs, while Bastani et al. [4], Wang et al. [29], Greenberg et al. [12], Choi et al. [6] develop risk-sensitive reinforcement learning algorithms. These works consider risk measures that fall into the class of coherent risk measures [2] that include the conditional value-at-risk and the entropic value-at-risk. Risk-sensitive objectives have also been considered in the stochastic control literature. Methods for risk-sensitive linear, quadratic, and Gaussian control [31] have been developed along with more recent methods for non-linear model predictive control [24]. We also refer to Wang and Chapman [30] for a comprehensive survey of this topic. However, all of these works stand in contrast with our framework where we want to minimize a risk measure of the maximum incurred loss instead of the cumulative loss.
@@
Related lines of research in robotics include hierarchical reinforcement learning [18, 7, 36, 16, 21] and task and motion planning (TAMP) [19] in which long-horizon control objectives are decomposed into subtasks that are carried out by motion planners or other low-level controllers trained with RL or imitation learning. Similar frameworks have been proposed for automatic generation of agentic workflows to automate software development tasks [25, 33, 15] that sequence LLM-agents to complete smaller subtasks. Agentic systems have also been proposed for the tasks of complex information retrieval [34] and scientific discovery [13]. These methods typically optimize the overall task success probability and do not consider risk minimization.
2 Risk Minimization Over Agent Graphs
@@
In this work, an agent is any machine learning model $f:X\to\mathfrak{D}(T\times Y)$ where $X$ is the input domain, $T$ is the set of computational traces capturing the set of agent behaviors, $Y$ is the output domain, and $\mathfrak{D}(T\times Y)$ is the set of probability distributions over $T\times Y$. For example, for an agent controlling a robot, $X$ is the set of initial states, $T$ is the set of trajectories of the robot in its environment, and $Y$ is the set of final states. In the case of a language model, $X$ is the set of prompts, $Y$ is the set of final responses, and $T$ is the set of intermediate tokens used to produce the final response. This can be the Chain-of-Thought reasoning traces, or interactions with a tool like retrieval from the web or running code. We consider a distribution over traces and outputs, as opposed to a single output, to enable modeling stochastic agent behaviors along with interactions with a stochastic environment. We only assume sampling access to the distribution over outputs and denote $(t,y)\sim f(x)$ for sampling a trace-output pair $(t,y)$ from the agent $f$ for the input $x$.
@@
We quantify the risk associated with an agent’s trace by a loss function $L:T\to\mathbb{R}$. This formulation allows us to decouple risk, which depends on the agent’s overall behavior, from task accuracy, which only depends on the final output. Thus, the loss function can encode desiderata of the agent’s behavior like safety, privacy, and fairness. For example, in the DroneNav environment from Example 1, the loss function quantifies how safe a control policy is on a trajectory of the drone.
@@
Observe that given a distribution over the inputs $\mathcal{D}(X)$, the agent $f$ induces distributions $\mathcal{D}_{f}(T)$ and $\mathcal{D}_{f}(Y)$ over the set of traces and outputs respectively. We can sample $t\sim\mathcal{D}_{f}(T)$ or $y\sim\mathcal{D}_{f}(Y)$ by first sampling $x\sim\mathcal{D}(X)$, then sampling $(t,y)\sim f(x)$, and dropping one of the two components. Evaluating the loss on the distribution of traces gives us a distribution over the losses the agent incurs. Since we are concerned with risk minimization, we focus on analyzing tail behaviors of the loss distribution. In particular, the quantity we are interested in optimizing is the value-at-risk at level $\alpha\in(0,1)$, denoted $\textnormal{VaR}_{\alpha}$. $\textnormal{VaR}_{\alpha}$ is a risk measure of the loss distribution that corresponds to the $(1-\alpha)$-quantile of the distribution. In other words, a $(1-\alpha)$ fraction of the traces will have a loss smaller than the $\textnormal{VaR}_{\alpha}$. To formally define it, let $Z_{f}\sim\mathcal{D}_{f}(T)$ be a random variable over the traces of the agent $f$. Then we can express
@@
which is the maximum of losses incurred by the sequence of trajectories $(t_{1},\dots,t_{m})$.
We now have all the ingredients to state our risk minimization objective over an agent graph (RMAG).
Objective (RMAG).
Let $G=(V,E,X,T,F,L,s,t,\mathcal{D}_{s})$ be an agent graph and let $\mathcal{P}\subseteq V^{*}$ be the set of directed paths from $s$ to $t$. For a given risk level $\alpha\in(0,1)$ we define the risk minimization objective as the following optimization problem
@@
3 Finding the path minimizes the value-at-risk
Algorithm BucketedVaR : Efficient algorithm for the RMAG objective